# A Branch-and-Cut-and-Price Algorithm for Cutting Stock and Related Problems

### v2

@@ Abstract
We present a branch-and-cut-and-price framework to solve Cutting Stock Problems with strong relaxations using Set Covering (Packing) Formulations, which are solved by column generation. The main contributions of this paper include an extended Ryan-Foster scheme, which allows us to use this powerful branching scheme even in non-binary problems by using a conflict propagation lemma; a fast column generation process based on a diversification strategy; custom primal heuristics, enabling us to find optimal solutions for several open instances; and a technique to use a smaller feasibility tolerance in floating-point linear programming solvers, combined with numerically safe methods to produce stronger and safer lower bounds. Additional performance-improving strategies include a technique that controls the height of the branch-and-bound tree; a variable selection algorithm based on branching history; a new set of dual inequalities; insights to obtain a lean model; and the subset-row inequalities. By employing this comprehensive framework, we overcame the current state-of-the-art concerning the following problems: Cutting Stock, Skiving Stock, Ordered Open-End Bin Packing, Class-Constrained Bin Packing, and Identical Parallel Machines Scheduling with Minimum Makespan. Additionally, a new challenging benchmark for Cutting Stock is introduced.
Keywords: Branch-and-Cut-and-Price; Set Covering Formulation; Variable Selection; Cutting Stock; Bin Packing; Scheduling.
@@ 1 Introduction
This paper introduces a branch-and-cut-and-price framework tailored to solve problems with strong relaxations using the SCF or the Set Packing Formulation (SPF). While our primary focus lies on addressing the CSP, we extend our framework to tackle the Skiving Stock Problem (SSP), Identical Parallel Machines Scheduling with Minimum Makespan (IPMS), Ordered Open-End Bin Packing Problem (OOEBPP), and Class-Constrained Bin Packing Problem (CCBPP). The main contributions of this paper are: an extended Ryan-Foster scheme, which allows us to use this powerful branching scheme even in non-binary problems by using a conflict propagation lemma; a fast column generation process based on a diversification strategy; custom primal heuristics, enabling us to find optimal solutions for several open instances; and a technique to use a smaller feasibility tolerance in Floating-point Linear Programming Solvers (FLPSs), combined with numerically safe methods to produce stronger and safer lower bounds. Additional performance-improving strategies include the splay operation, a technique that controls the height of the branch-and-bound (B&B) tree; a variable selection algorithm based on branching history; a new set of dual inequalities; insights to obtain a lean model; and the subset-row inequalities. By adopting this comprehensive framework, we overcame the current state-of-the-art in the five studied problems, demonstrating the framework’s extensibility and its potential applicability to a broader range of problems.
@@
The remaining sections of this paper are structured as follows. Section 2 outlines our branch-and-cut-and-price algorithm, detailing the components including our pricer, cutting planes, branching scheme, and an overview of other key features of our framework. Sections 3 and 4 introduce our conflict propagation mechanism and the method for using a smaller tolerance in FLPSs while ensuring numerical safety, respectively. Section 5 develops the column-and-row generation algorithms. The implementation details of the Relax-and-Fix primal heuristic, along with a variant proposed by us, are explained in Section 6. In Section 7, we describe our branching rule, the splay operation, our dual inequalities, the waste optimization, and a variable elimination by reduced cost method, where the latter two are aimed at achieving a leaner model. Furthermore, Section 8 presents the computational experiments conducted for the CSP, while Section 9 demonstrates how our framework can be adapted for the other four problems studied. Finally, Section 10 draws our conclusions and outlines potential avenues for future research.
2 Branch-and-Cut-and-Price Algorithm
@@
In the context of the CSP, a pattern $P$ represents a feasible way of cutting a stock roll to obtain the demanded items, satisfying the condition $\sum_{i=1}^{n}a_{i}^{P}w_{i}\leq W$, where $a_{i}^{P}$ denotes the number of copies of item $i$ in $P$. A pattern is considered valid if for every pair of items $\{i,j\}\in P$ they are not in a conflict, i.e., $\{i,j\}\notin E$. Thereafter, we will simply use the term pattern to refer to a valid pattern. A possible formulation for the CSP is the SCF proposed Gilmore Gomory [15], which we describe next:
@@
These inequalities are well-known for defining facets of the Set Partition Formulation  [6], which is also a valid formulation for the CSP. The presented inequalities are termed weak because they are limited to items with unitary demands. Note that we can derive the SCF by removing the less-than inequalities from set partition formulation, making the model (1)–(4) equally applicable to the CSP. SRI and its generalization, known as Rank 1 Chavátal-Gomory inequalities, have seen widespread use in recent decades due to their efficacy in practical problems, especially in BPP (see, e.g., Wei et al. [32], Baldacci et al. [1]) and VRP (see, e.g., Jepsen et al. [20], Pessoa et al. [27]).
@@
A B&B that uses a formulation based on CRG is referred to as branch-and-cut-and-price (B&C&P). In the following sections, we present our branching scheme in more details and provide an overview of our B&C&P framework, outlining additional features used to enhanced its performance.
@@ 2.2 Overview
Next, we present an overview of our framework, a branch-and-cut-and-price (B&C&P) algorithm based on SCF. Our framework’s success is based on primal heuristics, fast-increasing lower-bound techniques, and a fast relaxation solver. We outline the pseudocode of our framework in Algorithm 1.
@@ 13                   Remove nodes using the splay operation
Next, the B&C&P phase begins. At each node, we strengthen the RLM solution from the SCF using weak SRIs with three rows. Our algorithm incorporates custom heuristics, which are activated based on predefined thresholds. As custom heuristics previously used in literature are insufficient to solve the challenging instance class AI of CSP, we present a novel approach that combines Relax-and-Fix (RF) [3] and a variant proposed by us named Constrained Relax-and-Fix (CRF). In comparative evaluations, our custom heuristics demonstrate competitiveness with Gurobi’s general heuristics, as utilized by Lima et al. [22], the current state-of-the-art for CSP.
Subsequently, we adhere to a depth-first order in the B&C&P, prioritizing the left branch. Our branching scheme, the extended Ryan-Foster scheme that we propose, employs a rule based on past decisions and incorporates a conflict propagation strategy among items to break certain symmetries of the problem.
If the current node can be pruned by the numerically safe dual bound, we employ a new strategy termed splay operation, which drops off potentially unfruitful nodes from the B&C&P tree to reduce the execution time. We reevaluate the current node if this strategy results in the removal of at least one node. Otherwise, we proceed to explore the next unexplored right node. Furthermore, we terminate the algorithm as soon as we detect an optimal solution, which may already occur in Line 2.
Finally, we implement a lean RLM to enhance our algorithm’s efficiency by implementing two strategies to reduce its size. The first strategy involves limiting the maximum waste of a pattern while solving both the root node and subsequent B&C&P nodes. This technique was employed by Lima et al. [22], but it was applied only after the root node was solved (not during its solution). The second strategy entails a routine that relies on reduced costs to temporarily eliminate patterns that cannot belong to an optimal solution. These several techniques give a fast relaxation solver and a fast lower bound increase, enabling the efficient resolution of non-IRUP CSP instances, such as those in the ANI class.
@@ 4 Numeric Safety and Smaller Tolerance
Note that, if there is a pattern $P$ with $\overline{c}_{P}(\overline{\pi}^{\text{int}},\overline{\rho}^{\text{int}},K)<\frac{-K}{M}$, then $\overline{c}_{P}(\overline{\pi}^{\text{dim}},\overline{\rho}^{\text{dim}},1)<\frac{-1}{M}\leq-\epsilon$. Since all dual values were not increased in the diminished dual solution, it follows that $\overline{c}_{P}(\overline{\pi},\overline{\rho},1)\leq\overline{c}_{P}(\overline{\pi}^{\text{dim}},\overline{\rho}^{\text{dim}},1)$, and, thus, $P$ will be pivoted when added to RLM. Thus, our CRG process stops when ${\overline{c}_{P}(\overline{\pi}^{\text{int}},\overline{\rho}^{\text{int}},K)\geq\frac{-K}{M}}$ for all patterns $P\in\mathcal{P}$. At this point, we employ an alternative pricer, explained in Section 5.3, to determine a lower bound $B$ for the minimum reduced cost $\overline{c}(\overline{\pi}^{\text{int}},\overline{\rho}^{\text{int}},K)$, and we use $B$ to compute $z^{\text{safe}}$ (which, by the result of Baldacci et al. [1], is also a valid and safe bound).
@@
Concerning the second issue, if our algorithm detects it in $(\overline{\pi},\overline{\rho})$, we discard the current simplex basis and obtain a new one by solving RLM using the barrier method, which is faster than the simplex algorithm for solving from scratch. Using the barrier method, we never detected the second numerical issue on $(\overline{\pi},\overline{\rho})$. However, if it does arise, we could halt the CRG process and calculate a safe bound $z^{\text{safe}}$ with the current smallest reduced cost. Additionally, we have mitigated occurrences of the second numerical issue by setting Gurobi’s parameter Numeric Focus to 2, prompting the solver to exercise more caution during numerical manipulation. To detect if $(\overline{\pi},\overline{\rho})$ has the second issue, we assess it under $(\overline{\pi}^{\text{int}},\overline{\rho}^{\text{int}})$ by verifying if $\overline{c}_{P}(\overline{\pi}^{\text{int}},\overline{\rho}^{\text{int}},K)<\frac{-K}{M}$ for some pattern $P\in\overline{\mathcal{P}}$. Even though this could lead to false positives and false negatives, this is manageable. A false positive results in running the barrier method unnecessarily, while a false negative prevents the pricer from generating this column, thereby avoiding potential loops. Furthermore, for a false positive to occur, conversion errors from $(\overline{\pi},\overline{\rho})$ to $(\overline{\pi}^{\text{int}},\overline{\rho}^{\text{int}})$ must generate an accumulated error greater than $(\frac{1}{M}-\epsilon^{\prime})\cdot K\approx 640$ in the reduced cost of a pattern, which is not expected.
@@ 23      MultiplePatternGeneration($\overline{P},i,r$)
Finally, the alternative pricer mentioned in Section 4 for determining a lower bound $B$ for $\overline{c}(\overline{\pi}^{\text{int}},\overline{\rho}^{\text{int}},K)$ uses a heap-based approach. It replaces the recursive function’s arguments with labels $(\overline{P},i,r)$ and processes the label with the smallest lower bound (which no longer needs to be smaller than $\frac{-K}{M}$) that can improve $\overline{P}^{\text{inc}}$. Note that $B$ is the minimum of the reduced cost of $\overline{P}^{\text{inc}}$ and the lower bound of the unprocessed labels. For performance, we use an iteration limit $N^{R}_{\text{max}}=\frac{\overline{I}W}{50}$ and also halt if $B\geq-K\cdot 10^{-13}$. In this case, the difference between $z^{\text{safe}}$ obtained using $B$ and $\overline{c}(\overline{\pi}^{\text{int}},\overline{\rho}^{\text{int}},K)$ is negligible.
@@ 5.4 Generation of Subset-Row Inequalities
Proof. By definition, $S$ induces a violated clique cut if $\sum_{P\in\mathcal{P}}\lfloor\frac{a_{i}^{P}+a_{j}^{P}+a_{k}^{P}}{2}\rfloor\lambda_{P}^{*}>1$. Note that $\delta_{ij}+\delta_{jk}+\delta_{ki}=\sum_{P\in\mathcal{P}}(a_{i}^{P}a_{j}^{P}+a_{i}^{P}a_{k}^{P}+a_{j}^{P}a_{k}^{P})\lambda_{P}^{*}\geq\sum_{P\in\mathcal{P}}\lfloor\frac{a_{i}^{P}+a_{j}^{P}+a_{k}^{P}}{2}\rfloor\lambda_{P}^{*}>1$, since, if the coefficient on the right side is $1$, there are at least two items of the triple in $P$, ensuring the left side is at least $1$. Now, suppose $S$ is a cut such that $\delta_{ij}+\delta_{jk}+\delta_{ki}>1$ and only $\delta_{ij}>0$, without loss of generality. This implies that $\delta_{ij}>1$, which can only occur if items $i$ and $j$ are over-covered. Since we can convert any (fractional) solution with over-covered items into an exact-covered solution with the same value, $S$ cannot improve the lower bound.  $\blacksquare$
@@ 7.3 Dual Inequalities and Binary Pricing
The Dual Inequalities and Binary Pricing techniques discussed below aim to solve the root node without any SRIs, thus we omit $\overline{\rho}$ in dual solutions. Dual inequalities [4, 17] offer another promising strategy to address the slow convergence to solve the root node in CG-based formulations. These studies reveal that there exists a dual optimal solution $\overline{\pi}^{*}$ for LM, where the value $\overline{\pi}_{i}^{*}$ of each item $i$ is proportional to its size $w_{i}$. Consequently, stabilizing the dual values $\overline{\pi}$ can be done by adding constraints of the form ${\overline{\pi}_{i}\geq\sum_{j\in S}t_{j}\overline{\pi}_{j}}$ where ${w_{i}\geq\sum_{j\in S}t_{s}w_{j}}$, $S\subseteq\overline{I}\setminus{i}$, and $t\in\mathbb{Z}^{S}_{>0}$. Given the super-polynomial number of such inequalities, these authors often employ fixed-size subsets or dynamically separate them. Importantly, these inequalities do not compromise relaxation strength, and in cases where they yield an infeasible primal solution, conversion to a feasible one is straightforward.

### v3

# Solving Cutting Stock Problems via an Extended Ryan-Foster Branching Scheme and Fast Column Generation

@@ Abstract
We present a branch-cut-and-price framework to solve Cutting Stock Problems with strong relaxations using Set Covering (Packing) Formulations, which are solved by column generation. The main contributions of this paper include an extended Ryan-Foster scheme, which allows us to use this powerful branching scheme even in non-binary problems by using a conflict propagation lemma; a fast column generation process based on a diversification strategy; custom primal heuristics, enabling us to find optimal solutions for several open instances; and a technique to use a smaller feasibility tolerance in floating-point linear programming solvers, combined with numerically safe methods to produce stronger and safer lower bounds. Additional performance-improving strategies include a technique that controls the height of the branch-and-bound tree; a variable selection algorithm based on branching history; a new set of dual inequalities; insights to obtain a lean model; and the subset-row inequalities. By employing this comprehensive framework, we overcame the current state-of-the-art concerning the following problems: Cutting Stock, Skiving Stock, Ordered Open-End Bin Packing, Class-Constrained Bin Packing, and Identical Parallel Machines Scheduling with Minimum Makespan. Additionally, a new challenging benchmark for Cutting Stock is introduced.
Keywords: Branch-Cut-and-Price; Set Covering Formulation; Variable Selection; Cutting Stock; Bin Packing
@@ 1 Introduction
This paper introduces a branch-cut-and-price framework tailored to solve problems with strong relaxations using the SCF or the Set Packing Formulation (SPF). While our primary focus lies on addressing the CSP, we extend our framework to tackle the Skiving Stock Problem (SSP), Identical Parallel Machines Scheduling with Minimum Makespan (IPMS), Ordered Open-End Bin Packing Problem (OOEBPP), and Class-Constrained Bin Packing Problem (CCBPP). The main contributions of this paper are: an extended Ryan-Foster scheme, which allows us to use this powerful branching scheme even in non-binary problems by using a conflict propagation lemma; a fast column generation process based on a diversification strategy; custom primal heuristics, enabling us to find optimal solutions for several open instances; and a technique to use a smaller feasibility tolerance in Floating-point Linear Programming Solvers (FLPSs), combined with numerically safe methods to produce stronger and safer lower bounds. Additional performance-improving strategies include the splay operation, a technique that controls the height of the branch-and-bound (B&B) tree; a variable selection algorithm based on branching history; a new set of dual inequalities; insights to obtain a lean model; and the subset-row inequalities. By adopting this comprehensive framework, we overcame the current state-of-the-art in the five studied problems, demonstrating the framework’s extensibility and its potential applicability to a broader range of problems.
A preliminary version of this paper was published as part of an undergraduate research award competition of a Brazilian conference [25].
This preliminary version lacks the depth and detail of the current paper, as it is a brief report on the research conducted. Thus, the preliminary version has all proofs omitted, including the proof of the conflict propagation lemma, which is essential for our algorithm’s correctness. It also omits important details, such as how to implement the variable selection rule based on branching history, for example.
This paper also presents new contributions that are not included in the preliminary version. It introduces a new set of dual inequalities and a numerically safe method to obtain lower bounds. It also extends the results of the preliminary version to three additional problems (IPMS, OOEBPP, and CCBPP), whereas the preliminary version only considers the CSP and SSP. Furthermore, it discusses the necessary adaptations to address these problems and presents detailed computational results for all five problems. Finally, it introduces a new benchmark for the CSP, which may be of great interest for future research and is not included in the preliminary version.
@@
The remaining sections of this paper are structured as follows. Section 2 outlines our branch-cut-and-price algorithm, detailing the components including our pricer, cutting planes, branching scheme, and an overview of other key features of our framework. Sections 3 and 4 introduce our conflict propagation mechanism and the method for using a smaller tolerance in FLPSs while ensuring numerical safety, respectively. Section 5 develops the column-and-row generation algorithms. The implementation details of the Relax-and-Fix primal heuristic, along with a variant proposed by us, are explained in Section 6. In Section 7, we describe our branching rule, the splay operation, our dual inequalities, the waste optimization, and a variable elimination by reduced cost method, where the latter two are aimed at achieving a leaner model. Furthermore, Section 8 presents the computational experiments conducted for the CSP, while Section 9 demonstrates how our framework can be adapted for the other four problems studied. Finally, Section 10 draws our conclusions and outlines potential avenues for future research.
2 Branch-Cut-and-Price Algorithm
@@
In the context of the CSP, a pattern $P$ represents a feasible way of cutting a stock roll to obtain the demanded items, satisfying the condition $\sum_{i=1}^{\overline{n}}a_{i}^{P}w_{i}\leq W$, where $a_{i}^{P}$ denotes the number of copies of item $i$ in $P$. A pattern is considered valid if for every pair of items $\{i,j\}\in P$ they are not in a conflict, i.e., $\{i,j\}\notin E$. Thereafter, we will simply use the term pattern to refer to a valid pattern. A possible formulation for the CSP is the SCF proposed Gilmore Gomory [15], which we describe next:
@@
These inequalities are well-known for defining facets of the Set Partition Formulation  [6], which is also a valid formulation for the CSP. The presented inequalities are termed weak because they are limited to items with unitary demands. SRI and its generalization, known as Rank 1 Chavátal-Gomory inequalities, have seen widespread use in recent decades due to their efficacy in practical problems, especially in BPP (see, e.g., Wei et al. [33], Baldacci et al. [1]) and VRP (see, e.g., Jepsen et al. [20], Pessoa et al. [28]).
@@
A B&B that uses a formulation based on CRG is referred to as branch-cut-and-price (BC&P). In the following sections, we present our branching scheme in more details and provide an overview of our BC&P framework, outlining additional features used to enhanced its performance.
@@ 2.2 Overview
Next, we present an overview of our framework, a branch-cut-and-price (BC&P) algorithm based on SCF. Our framework’s success is based on primal heuristics, fast-increasing lower-bound techniques, and a fast relaxation solver. We outline the pseudocode of our framework in Algorithm 1.
@@ 13                   Remove nodes using the splay operation
Next, the BC&P phase begins. At each node, we strengthen the RLM solution from the SCF using weak SRIs with three rows. Our algorithm incorporates custom heuristics, which are activated based on predefined thresholds. As custom heuristics previously used in literature are insufficient to solve the challenging instance class AI of CSP, we present a novel approach that combines Relax-and-Fix (RF) [3] and a variant proposed by us named Constrained Relax-and-Fix (CRF). In comparative evaluations, our custom heuristics demonstrate competitiveness with Gurobi’s general heuristics, as utilized by Lima et al. [22], the current state-of-the-art for CSP.
Subsequently, we adhere to a depth-first order in the BC&P, prioritizing the left branch. Our branching scheme, the extended Ryan-Foster scheme that we propose, employs a rule based on past decisions and incorporates a conflict propagation strategy among items to break certain symmetries of the problem.
If the current node can be pruned by the numerically safe dual bound, we employ a new strategy termed splay operation, which drops off potentially unfruitful nodes from the BC&P tree to reduce the execution time. We reevaluate the current node if this strategy results in the removal of at least one node. Otherwise, we proceed to explore the next unexplored right node. Furthermore, we terminate the algorithm as soon as we detect an optimal solution, which may already occur in Line 2.
Finally, we implement a lean RLM to enhance our algorithm’s efficiency by implementing two strategies to reduce its size. The first strategy involves limiting the maximum waste of a pattern while solving both the root node and subsequent BC&P nodes. This technique was employed by Lima et al. [22], but it was applied only after the root node was solved (not during its solution). The second strategy entails a routine that relies on reduced costs to temporarily eliminate patterns that cannot belong to an optimal solution. These several techniques give a fast relaxation solver and a fast lower bound increase, enabling the efficient resolution of non-IRUP CSP instances, such as those in the ANI class.
@@ 4 Numeric Safety and Smaller Tolerance
Note that, if there is a pattern $P$ with $\overline{c}_{P}(\overline{\pi}^{\text{int}},\overline{\rho}^{\text{int}},K)<\frac{-K}{M}$, then $\overline{c}_{P}(\overline{\pi}^{\text{dim}},\overline{\rho}^{\text{dim}},1)<\frac{-1}{M}\leq-\epsilon$. Since all dual values were not increased in the diminished dual solution, it follows that $\overline{c}_{P}(\overline{\pi},\overline{\rho},1)\leq\overline{c}_{P}(\overline{\pi}^{\text{dim}},\overline{\rho}^{\text{dim}},1)$, and, thus, $P$ will be pivoted when added to RLM. Therefore, our CRG process stops when ${\overline{c}_{P}(\overline{\pi}^{\text{int}},\overline{\rho}^{\text{int}},K)\geq\frac{-K}{M}}$ for all patterns $P\in\mathcal{P}$. At this point, we employ an alternative pricer, explained in Section 5.3, to determine a lower bound $B$ for the minimum reduced cost $\overline{c}(\overline{\pi}^{\text{int}},\overline{\rho}^{\text{int}},K)$, and we use $B$ to compute $z^{\text{safe}}$ (which, by the result of Baldacci et al. [1], is also a valid and safe bound).
@@
Concerning the second issue, if our algorithm detects it in $(\overline{\pi},\overline{\rho})$, we discard the current simplex basis and obtain a new one by solving RLM using the barrier method, which is faster than the simplex algorithm for solving from scratch. Using the barrier method, we never detected the second numerical issue on $(\overline{\pi},\overline{\rho})$. However, if it does arise, we could halt the CRG process and calculate a safe bound $z^{\text{safe}}$ with the current smallest reduced cost. Additionally, we have mitigated occurrences of the second numerical issue by setting Gurobi’s parameter Numeric Focus to 2, prompting the solver to exercise more caution during numerical manipulation. To detect if $(\overline{\pi},\overline{\rho})$ has the second issue, we assess it under $(\overline{\pi}^{\text{int}},\overline{\rho}^{\text{int}})$ by verifying if $\overline{c}_{P}(\overline{\pi}^{\text{int}},\overline{\rho}^{\text{int}},K)<\frac{-K}{M}$ for some pattern $P\in\overline{\mathcal{P}}$, i.e., for the patterns in RLM. Even though this could lead to false positives and false negatives, this is manageable. A false positive results in running the barrier method unnecessarily, while a false negative prevents the pricer from generating this column, thereby avoiding potential loops. Furthermore, for a false positive to occur, conversion errors from $(\overline{\pi},\overline{\rho})$ to $(\overline{\pi}^{\text{int}},\overline{\rho}^{\text{int}})$ must generate an accumulated error greater than $(\frac{1}{M}-\epsilon^{\prime})\cdot K\approx 640$ in the reduced cost of a pattern, which is not expected.
@@ 23      MultiplePatternGeneration($\overline{P},i,r$)
Finally, the alternative pricer mentioned in Section 4 for determining a lower bound $B$ for $\overline{c}(\overline{\pi}^{\text{int}},\overline{\rho}^{\text{int}},K)$ uses a heap-based approach. It replaces the recursive function’s arguments with labels $(\overline{P},i,r)$ and processes the label with the smallest lower bound (which no longer needs to be smaller than $\frac{-K}{M}$) that can improve $\overline{P}^{\text{inc}}$. Note that $B$ is the minimum of the reduced cost of $\overline{P}^{\text{inc}}$ and the lower bound of the unprocessed labels. For performance, we use an iteration limit $N^{R}_{\text{max}}=\frac{|\overline{I}|W}{50}$ and also halt if $B\geq-K\cdot 10^{-13}$. In this case, the difference between $z^{\text{safe}}$ obtained using $B$ and $\overline{c}(\overline{\pi}^{\text{int}},\overline{\rho}^{\text{int}},K)$ is negligible.
@@ 5.4 Generation of Subset-Row Inequalities
Proof. By definition, $S$ induces a violated clique cut if $\sum_{P\in\mathcal{P}}\lfloor\frac{a_{i}^{P}+a_{j}^{P}+a_{k}^{P}}{2}\rfloor\lambda_{P}^{*}>1$. Note that $\delta_{ij}+\delta_{jk}+\delta_{ki}=\sum_{P\in\mathcal{P}}(a_{i}^{P}a_{j}^{P}+a_{i}^{P}a_{k}^{P}+a_{j}^{P}a_{k}^{P})\lambda_{P}^{*}\geq\sum_{P\in\mathcal{P}}\lfloor\frac{a_{i}^{P}+a_{j}^{P}+a_{k}^{P}}{2}\rfloor\lambda_{P}^{*}>1$, since, if the coefficient on the right side is $1$, there are at least two items of the triple in $P$, ensuring the left side is at least $1$. Now, suppose $S$ is a cut such that $\delta_{ij}+\delta_{jk}+\delta_{ki}>1$ and only $\delta_{ij}>0$, without loss of generality. This implies that $\delta_{ij}>1$, which can only occur if $i$ and $j$ are over-covered. Since we can convert any (fractional) solution with over-covered items into an exact-covered solution with the same value, $S$ cannot improve the lower bound.  $\blacksquare$
@@ 7.3 Dual Inequalities and Binary Pricing
The Dual Inequalities and Binary Pricing techniques discussed below aim to solve the root node without any SRIs, thus we omit $\overline{\rho}$ in dual solutions. Dual inequalities [4, 17] offer another promising strategy to address the slow convergence to solve the root node in CG-based formulations. These studies reveal that there exists a dual optimal solution $\overline{\pi}^{*}$ for LM, where the value $\overline{\pi}_{i}^{*}$ of each item $i$ is proportional to its size $w_{i}$. Consequently, stabilizing the dual values $\overline{\pi}$ can be done by adding constraints of the form ${\overline{\pi}_{i}\geq\sum_{j\in S}t_{j}\overline{\pi}_{j}}$ where ${w_{i}\geq\sum_{j\in S}t_{s}w_{j}}$, $S\subseteq\overline{I}\setminus{i}$, and $t\in\mathbb{Z}^{S}_{>0}$. Given the superpolynomial number of such inequalities, these authors often employ fixed-size subsets or dynamically separate them. Importantly, these inequalities do not compromise relaxation strength, and in cases where they yield an infeasible primal solution, conversion to a feasible one is straightforward.